A food processor produces a sound level of 87dB at a distance of 1.0 m. How many watts per square metre does this sound level correspond to?
Answer #1

formula for sound intensity level is

beta = 10 dB log ( I/Io)

87 dB = 10 dB log ( I/Io)

I = Io * 10^87/10 = 1* 10^-12 (10^8.7) = 5.0* 10^-4 W/m^2